# Convergence of patterned matrices with random walk entries

### v1

@@ Abstract
It is well known that the Brownian motion on the real line can be obtained as a weak limit of a suitably scaled continuous-time random walk (CTRW). We investigate the convergence of certain patterned random matrices whose entries are independent CTRWs of various types. In a non-commutative probability framework, we use these high dimensional matrices to derive approximations of the free Brownian motion. Furthermore, we introduce and analyze a random time-changed version of the free Brownian motion driven by an inverse stable subordinator. An approximation of this process is obtained using a random matrix whose entries consist of continuous-time randomly stopped random walks. Moreover, it is shown that the empirical spectral distributions of such matrices have longer tails. Additionally, in a specific case, we use the explicit eigenvalue expressions of these matrices to obtain weak approximations of the standard Brownian motion and a time-changed variant of it.
@@
Keywords: Patterned matrix, Wigner matrix, fractional Poisson process, semi-circle law, Brownian motion, free Brownian motion, inverse stable subordinator.
@@ 1 Introduction
Let $\{X_{k}\}_{k\geq 1}$ be independent and identically distributed (iid) random variables (rvs) with mean zero and variance one, and let $\{S_{n}:=\sum_{k=1}^{n}X_{k}\}_{n\geq 0}$ with $S_{0}=0$. It is well known that the scaled continuous-time random walk (CTRW), $\{n^{-1/2}S_{[nt]}\}_{t\geq 0}$, $n\geq 1$, where $\cdot$] is the greatest integer function, converges weakly to a (standard) Brownian motion (SBM) $\{W(t)\}_{t\geq 0}$, as $n\rightarrow\infty$. This conclusion remains true if $\{X_{k}\}$ are not necessarily identically distributed but are independent with uniformly bounded moments of all orders. The SBM is a Gaussian process with independent and stationary increments, $W(0)=0$ and for all $0\leq s\leq t<\infty$, W(t)W(s)W(t)-W(s) is a Gaussian variable with mean zero and variance tst-s.
Let $\{N(t)\}_{t\geq 0}$ be a renewal process with iid inter-event times $\{J_{k}\}_{k\geq 1}$, independent of $\{X_{k}\}_{k\geq 1}$. Then $\{S_{N(t)}\}_{t\geq 0}$ is a time-changed random walk. If $\{N(t)\}_{t\geq 0}$ is a Poisson process with rate one and compounding rvs $\{X_{k}\}_{k\geq 1}$ are standard Gaussian then the scaled CTRW, $\{n^{-1/2}S_{N(nt)}\}_{t\geq 0}$, $n\geq 1$, also converges weakly in an appropriate Skorokhod topology to an SBM (see [26]). Moreover, if $\{J_{k}\}_{k\geq 1}$ belong to the domain of attraction of a stable law with index $\alpha\in(0,1)$ and $\{X_{k}\}_{k\geq 1}$ belong to the domain of attraction of a standard Gaussian law, then the scaled CTRW, $\{n^{-\alpha/2}S_{N(nt)}\}_{t\geq 0}$, $n\geq 1$, $0<\alpha<1$, weakly converges in an appropriate Skorokhod topology to a time-changed Brownian motion (see [23, 26]), where the time changing component is an inverse stable subordinator (for definition see Section 2.4).
In non-commutative probability, the analogues of the Gaussian distribution and the Brownian motion are respectively the semi-circle distribution (also known as the free Gaussian distribution) and the (standard) free Brownian motion (FBM), say $\{B(t)\}_{t\geq 0}$ (see Section 2.2) for which the increments are freely independent and the moments of $B(t)$ coincides with the moments of the semi-circle distribution with variance $t$. At the same time, it is known that the algebraic limit variable of scaled Wigner matrices of increasing dimensions has the semi-circle distribution, and that independent Wigner matrices are almost surely (a.s.) asymptotically free.
@@
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a classical probability space. Let $\mathcal{A}_{\mathbb{P}}=\cap_{1\leq p<\infty}L^{p}(\mathbb{P})$ be the algebra of rvs with all moments finite and let $\mathcal{M}_{n}(\mathcal{A}_{\mathbb{P}})$ be the algebra of $n\times n$ complex random matrices with entries from $\mathcal{A}_{\mathbb{P}}$. Then, the pair $(\mathcal{M}_{n}(\mathcal{A}_{\mathbb{P}}),\tau_{n})$ with $\tau_{n}=n^{-1}\mathbb{E}^{\mathbb{P}}\mathrm{Trace}(\cdot)$ or $\tau_{n}=n^{-1}\mathrm{Trace}(\cdot)$ forms a non-commutative probability space (see Section 2.1 for a formal definition). Let $\{A_{n}(t)\}_{t\geq 0}$ be a collection from $\mathcal{M}_{n}(\mathcal{A}_{\mathbb{P}})$. Then, it is said to be algebraically convergent to a collection $\{A(t)\}_{t\geq 0}$ in finite dimensional sense if for any $p\in\mathbb{N}$ and distinct $t_{1},\dots,t_{p}\in[0,\infty)$, the matrices $A_{n}(t_{1}),\dots,A_{n}(t_{p})$ jointly converge to variables $A(t_{1}),\dots,A(t_{p})$ (see Section 2.1.2 for details).
Motivated by the above developments, we consider some sequences of patterned random matrices whose entries evolve as independent CTRWs. Let $(S_{i,j,n})_{1\leq i,j\leq n}$, $n\geq 1$ be independent random walks whose steps are independent with mean 0, variance $1$ and all moments uniformly bounded. For $d\in\{1,2\}$, let $L_{n}:\{1,2,\dots,n\}^{2}\rightarrow\mathbb{Z}^{d}$ be link functions, with $L_{n+1}(i,j)=L_{n}(i,j)$ for all $1\leq i,j\leq n$. By abuse of notation, we denote these functions by $L$. For a suitable symmetric $L$ (that is, $L(i,j)=L(j,i)$ for all $i,j$), these yield symmetric random matrices $\{A_{n}(t)=(S_{L(i,j),[nt]})_{1\leq i,j\leq n}\}_{t\geq 0}$ in the NCP $(\mathcal{M}_{n}(\mathcal{A}_{\mathbb{P}}),\tau_{n})$. We also consider a time-changed version of this process $\{A_{n}^{\alpha}(t)=(S_{L(i,j),N^{\alpha}(t)})_{1\leq i,j\leq n}\}_{t\geq 0}$, $0<\alpha\leq 1$, where the time-changing component $\{N^{\alpha}(t)\}_{t\geq 0}$ is an independent fractional Poisson process (see Section 2.4 for the definition).
@@
After recalling in Section 2, some preliminaries that we need, in Sections 3.1 and 3.2, we establish respectively, the marginal and finite dimensional joint algebraic convergence as defined above (that is, convergence of the expected average trace of any polynomial) of $\{n^{-1}A_{n}(t)\}_{t\geq 0}$ for suitable links $L$. These algebraic convergence results also yield the convergence of the empirical spectral distribution of symmetric polynomials in independent copies of these matrices, evaluated at different time points.
@@
In particular, for the choice of $L(i,j)=(\max(i,j),\min(i,j))$, $\{n^{-1}A_{n}(t)\}_{t\geq 0}$ converges algebraically to a FBM $\{B(t)\}_{t\geq 0}$. The empirical spectral distributions of any symmetric polynomial in these matrices converges to the distributions determined by the moments of the corresponding (self-adjoint) polynomial of $\{B(t)\}_{t\geq 0}$. The choice of L(i,j)=n/2|n/2|ij||L(i,j)=n/2-|n/2-|i-j||, $1\leq i,j\leq n$, yields a Symmetric Circulant matrix. In this case, the matrices are commutative and $\{n^{-1}A_{n}(t)\}_{t\geq 0}$ converge algebraically to $\{W(t)\}_{t\geq 0}$, whose finite dimensional moments agree with those of an SBM.
@@
In Section 3.3, with the link function L(i,j)=n/2|n/2|ij||L(i,j)=n/2-|n/2-|i-j||, $1\leq i,j\leq n$, we consider the empirical process, $Y_{n}(t)=n^{-1}\sum_{r=1}^{n}\lambda_{r,n}(t)\mathbb{I}(U_{n}=r)$ where $\lambda_{r,n}(t),1\leq r\leq n$ are the eigenvalues of $A_{n}(t)$ and $U_{n}$ is a uniform random variable independent of all other variables. We show that $Y_{n}(\cdot)$ converges weakly to an SBM.
@@
In Section 4 we introduce the time-changed FBM. In Section 4.1, we consider patterned random matrices with stopped (via fractional Poisson process) random walk entries. we establish the finite dimensional algebraic convergence of the matrices $n^{-(1+\alpha)/2}A_{n}^{\alpha}(t)_{t\geq 0}$. In particular, for the choice $L(i,j)=(\max(i,j),\min(i,j))$ the limit is $\{B(L^{\alpha}(t))\}_{t\geq 0}$, where $\{L^{\alpha}(t)\}_{t\geq 0}$ is an inverse stable subordinator independent of $B$. The moments of $B(L^{\alpha}(t))$ identify a unique probability distribution. Simulations (see Figure 1) show that its density function has a longer tail compared to the semi-circle density with variance $t$, and as $\alpha\to 1$, the value of $\alpha$ increases, this density approaches the latter. In Section 4.2 we consider the link function L(i,j)=n/2|n/2|ij||L(i,j)=n/2-|n/2-|i-j||, $1\leq i,j\leq n$, and prove results similar to those of Section 3.3, the marginal limit in this case being $\{W(L^{\alpha}(t))\}_{t\geq 0}$, where $\{L^{\alpha}(t)\}_{t\geq 0}$ is an inverse stable subordinator independent of $W(t)$.
@@ 2.1 Non-commutative probability and (free) independence
A non-commutative probability space (NCP) is a pair $(\mathcal{A},\tau)$ with unital algebra $\mathcal{A}$ over $\mathbb{C}$ and a linear functional $\tau:\mathcal{A}\rightarrow\mathbb{C}$ such that $\tau(1_{\mathcal{A}})=1$, where $1_{\mathcal{A}}$ denotes the identity in $\mathcal{A}$. The state $\tau$ is called tracial if $\tau(ab)=\tau(ba)$ for all $a,b$ in $\mathcal{A}$. If $\mathcal{A}$ is a *-algebra, that is, there is an involution or *-operator $\mathcal{A}\rightarrow\mathcal{A}$. A state $\tau$ on $\mathcal{A}$ is called positive if $\tau(a^{*}a)\geq 0$ for all $a\in\mathcal{A}$. The pair $(\mathcal{A},\tau)$ with *-algebra $\mathcal{A}$ and positive state $\tau$ is called *-probability space.
@@ 2.1.2 Joint distribution and asymptotic freeness
2.2 Free Brownian motion
(i) $B(0)=0$;
2.3 Patterned random matrices and link function
Note that the notion of convergence as defined in Section 2.1.2 applies to random matrices with the choice of state being $\mathbb{E}\text{Trace}$. A related concept of convergence is that of the spectral distribution. Let $A_{n}$ be an $n\times n$ symmetric random matrix whose entries are rvs on a probability space $(\Omega,\mathcal{F},\mathbb{P})$, and its eigenvalues are $\lambda_{1},\dots,\lambda_{n}$. The empirical spectral distribution (ESD) of $A_{n}$ is defined as follows:
Wigner matrix is one of the most important random matrices, defined as $W_{n}:=((x_{ij}))_{n\times n}$ where $x_{ij}=x_{ji}$ and they are iid with mean 0 and variance $1$. The ESD of $n^{-1/2}W_{n}$ converges a.s. to the standard semi-circle law (see [33]). Independent scaled Wigner matrices also converge algebraically (when all moments are uniformly bounded) and are asymptotically free, the limits being free semi-circle variables. See [1] or [9].
A class of patterned random matrices that includes many important matrices has been studied in a unified way by [10]. We introduce the basic ideas of this approach now. For $d\in\{1,2\}$, let us consider a sequence $\{L_{n}\}_{n\geq 1}$ of functions defined as $L_{n}:\{1,2,\dots,n\}^{2}\rightarrow\mathbb{Z}^{d}$ with $L_{n+1}(i,j)=L_{n}(i,j)$. By using a common notation $L\coloneqq L_{n}$, it is called the link function. Now, by abuse of notation, $\mathbb{N}^{2}$ is a common domain of the sequence $\{L_{n}\}$.
Let $\{a_{i},\ i\geq 1\}$ and $\{a_{i,j},\ i\geq 1,\ j\geq 1\}$ be sequence and bi-sequence of rvs. For $n\geq 1$, a $n\times n$ random matrix $A_{n}$ with link function $L$ and input sequence $\{a_{i},\ i\geq 1\}$ or $\{a_{i,j},\ i\geq 1,\ j\geq 1\}$ is defined as $A_{n}\coloneqq((a_{L(i,j)}))_{1\leq i,j\leq n}$. The following are examples of link functions for some well known random matrices:
(i) $L(i,j)=(\max\{i,j\},\min\{i,j\})$: Wigner matrix;
(ii) L(i,j)=|ij|L(i,j)=|i-j|: (symmetric) Toepliz matrix;
(iii) L(i,j)=i+j2L(i,j)=i+j-2: (symmetric) Hankel matrix;
(iv) L(i,j)=i+j2modnL(i,j)=i+j-2\ $\text${mod}\ n, $1\leq i,j\leq n$: Reverse Circulant matrix;
(v) L(i,j)=n/2|n/2|ij||L(i,j)=n/2-|n/2-|i-j||, $1\leq i,j\leq n$: Symmetric Circulant matrix.
For several convergence results for patterned random matrices, both for the ESD and for joint algebraic convergence, see [9, 10, 12] and references therein. The following property that all the above link functions satisfy plays a crucial role:
Property B
$\Delta L\coloneqq\sup_{n}\sup_{r\in\mathbb{Z}^{d}_{+}}\sup_{1\leq i\leq n}\#\{j:1\leq j\leq n,\ L(i,j)=r\}<\infty$.
From (2.1), the $l$th moment associated to the ESD of $A_{n}$ is $\mu^{(l)}_{F_{A_{n}}}=n^{-1}\mathrm{Trace}(A_{n}^{l})$, $l\geq 1$, and the $l$th moment associated to its EESD is given by the trace-moment formula:
$\mathbb{E}\mu_{F_{A_{n}}}^{(l)}=n^{-1}\mathbb{E}\mathrm{Trace}(A_{n}^{l})=n^{-1}\sum_{1\leq j_{1},\dots,j_{l}\leq n}\mathbb{E}a_{L(j_{1},j_{2})}\dots a_{L(j_{l},j_{1})},\ l\geq 1.$
Let us consider a non random sequence $\{\mu^{(l)}\}_{l\geq 1}$ such that
(M1) For all $l\geq 1$, $\mathbb{E}\mu^{(l)}_{F_{A_{n}}}\rightarrow\mu^{(l)}$ as $n\rightarrow\infty$;
(M2) $\mathbb{V}\mathrm{ar}(\mu^{(l)}_{F_{A_{n}}})\rightarrow 0$ as $n\rightarrow\infty$.
If the sequence $\{\mu^{(l)}\}_{l\geq 1}$ satisfies Carleman’s condition, $\sum_{l=1}^{\infty}(\mu^{(2l)})^{-1/2l}=\infty$, then $\{\mu^{(l)}\}_{l\geq 1}$ uniquely determines a probability distribution, and in that case the ESD converges in probability to this distribution. The convergence can be strengthened to a.s. sense if (M2) is replaced by the following fourth moment condition:
(M4) for all $n\geq 1$, $\sum_{l=1}^{\infty}\mathbb{E}\big(\mu^{(l)}_{F_{A_{n}}}-\mathbb{E}\mu^{(l)}_{F_{A_{n}}}\big)^{4}<\infty$.
This route has been taken for patterned random matrices. The key step is to verify (M1). We outline how this is done using the trace-moment formula. For $l\geq 1$ and $n\geq 1$, a function $\pi:\{0,1,\dots,l\}\rightarrow\{1,2,\dots,n\}$ with $\pi(0)=\pi(l)$ is called a circuit of length $l\geq 1$. Thus, with $A_{\pi}=a_{L(\pi(0),\pi(1))}\dots a_{L(\pi(l-1),\pi(l))}$,
$\mathbb{E}\mu_{F_{A_{n}}}^{(l)}=n^{-1}\sum_{\pi:\,\text{$\pi$ is circuit of length $l$}}\mathbb{E}A_{\pi}.$
A circuit $\pi$ is called matched if all the $L$-values $L(\pi(j-1),\pi(j))$ are repeated. It is called pair-matched if each value $L(\pi(j-1),\pi(j))$ appears exactly twice. We group these circuits into equivalence classes: circuits $\pi_{1}$ and $\pi_{2}$ belong to the same class, denoted as $\pi_{1}\sim\pi_{2}$ if $L(\pi_{1}(j-1),\pi_{1}(j))=L(\pi_{1}(i-1),\pi_{1}(i))$ if and only if $L(\pi_{2}(j-1),\pi_{2}(j))=L(\pi_{2}(i-1),\pi_{2}(i))$ for all $1\leq i,j\leq l$. These equivalence classes can be indexed by partitions of $\{1,\dots,l\}$, which are represented by words $w$ of length $l$ in which letters appear in alphabetical order. Let $w(i)$ denote the $i$th letter of $w$ and $\Pi(w)$ be the equivalence class associated with word $w$. Then,
$\Pi(w)\coloneqq\{\pi:w(i)=w(j)\ \text{if and only if}\ L(\pi(j-1),\pi(j))=L(\pi(i-1),\pi(i))\}.$
Hence,
$\mathbb{E}\mu_{F_{A_{n}}}^{(l)}=\sum_{w:\ \text{$w$ is a word of length $l$}}n^{-1}\sum_{\pi\in\Pi(w)}\mathbb{E}A_{\pi}.$ (2.2)
A word $w$ is called matched if each letter repeats, and it is called pair-matched if each letter repeats exactly twice. Note that a pair-matched word is always of even length. Let $\mathcal{P}_{2}(2m)$ denote the set of pair-matched word of length $2m$, $m\geq 1$. It is then shown that the inner sum has a limit for each $w$ (and is 0 if $w\notin\mathcal{P}_{2}(2m)$ for any $m$). This then gives the convergence of the ESD and also implies algebraic convergence. For details, see [9].
2.3.1 Joint convergence of independent random matrices
The same ideas can be used for establishing joint algebraic convergence. Let $A_{r,n}$, $r=1,2,\dots,p$ be $n\times n$ independent random matrices with common link function $L$. The $(i,j)$th entry of $A_{r,n}$ will be denoted by $A_{r,n}(L(i,j))$. Note that a typical monomial $g$ in terms of $\{A_{r,n}\}_{1\leq r\leq p}$ has the following form: $g(\{A_{r,n}\}_{1\leq r\leq p})=A_{r_{1},n}\dots A_{r_{l},n}$, where $r_{j}\in\{1,2,\dots,p\}$ for each $j=1,2,\dots,l$. The collection $\{A_{r,n}\}_{1\leq r\leq p}$ in $(\mathcal{M}_{n}(\mathbb{C}),\tau_{n}=n^{-1}\mathbb{E}\mathrm{Trace})$ converges jointly if for every fixed monomial $g$, the following converges:
$\tau_{n}(g)=n^{-1}\sum_{\pi:\,\text{$\pi$ is a circuit of length $m$}}\mathbb{E}A_{r_{1},n}(L(\pi(0),\pi(1)))\dots A_{r_{l}n}(L(\pi(l-1),\pi(l))).$
It turns out that monomials $g$ with odd number of factors always yield the limit 0 and hence without loss we assume that the order of $g$ is $2m$ for some integer $m$. A circuit $\pi$ of length $l$ is called index-matched if the value $(r_{j},L(\pi(j-1),\pi(j)))$ repeats for each $j=1,2,\dots,l$. Note that circuits are index-matched with reference to the monomial $g$. Let $I=\{\pi:\ \pi\ \text{is a index-matched circuit}\}$. Consider a equivalence class on $I$ where for any two circuits $\pi_{1}$ and $\pi_{2}$, we have $\pi_{1}\sim\pi_{2}$ if for any $1\leq i,j\leq l$ such that $r_{i}=r_{j}$, $A_{r_{i},n}(L(\pi_{1}(i-1),\pi_{1}(i)))=A_{r_{j},n}(L(\pi_{1}(j-1),\pi_{1}(j)))$ if and only if $A_{r_{i},n}(L(\pi_{2}(i-1),\pi_{2}(i)))=A_{r_{j},n}(L(\pi_{2}(j-1),\pi_{2}(j)))$. Any such equivalence class can be indexed by a partition of $\{1,2,\dots,l\}$ which further induces an indexed word $w$ of length $l$, where the indexed letters appear in alphabetical order. An indexed word is called pair-matched if every indexed letter in it appears exactly twice. Only such letters remain relevant in the limit. Let $w_{r_{j}}(j)$ denote the $j$th letter of indexed matched word $w$ and $\Pi_{g}(w)$ denote the class of circuits corresponding to $w$ and monomial $g$. Then,
$\Pi_{g}(w)=\{\pi:\ w_{r_{i}}(i)=w_{r_{j}}(j)\ \text{iff}\ A_{r_{i},n}(L(\pi(i-1),\pi(i)))=A_{r_{j},n}(L(\pi(j-1),\pi(j)))\}.$
Using the results for a single matrix, this would show the joint convergence of the iid copies of the pattern matrices described above. For further details, see [6, 9, 10, 11], and references therein.

### v2

# Patterned matrices with random walk entries

@@ Abstract
It is well known that the weak limit of a suitably scaled continuous-time random walk (CTRW) is the Brownian motion. We investigate the convergence of certain patterned random matrices whose entries are independent CTRWs and their time-changed versions, in a non-commutative probability framework. For the Wigner link function, the limits are free Brownian motion and its time-changed version driven by an inverse stable subordinator. For the symmetric circulant and the circulant with CTRW entries, we use their explicit eigenvalue expressions to define some empirical processes that converge weakly to a Brownian motion and a complex Brownian motion, respectively. For matrices with iid entries, and for elliptic matrices, the algebraic limits are equal in *-distribution to processes whose marginals are circular and elliptic variables, respectively. A random time-changed variant of these results is also established.
@@
Keywords: Patterned matrix, Wigner matrix, circulant matrix, elliptic matrix, fractional Poisson process, semi-circle law, Brownian motion, free Brownian motion, inverse stable subordinator.
@@ 1 Introduction
Let $\{X_{k}\}_{k\geq 1}$ be independent and identically distributed (iid) random variables (rvs) with mean zero and variance one, and let $\{S_{n}:=\sum_{k=1}^{n}X_{k}\}_{n\geq 0}$ with $S_{0}=0$. It is well known that the scaled continuous-time random walk (CTRW), $\{n^{-1/2}S_{[nt]}\}_{t\geq 0}$, $n\geq 1$, where $\cdot$] is the greatest integer function, converges weakly to a (standard) Brownian motion (BM) $\{W(t)\}_{t\geq 0}$, as $n\rightarrow\infty$. This conclusion remains true if $\{X_{k}\}$ are not necessarily identically distributed but are independent with uniformly bounded moments of all orders. The BM is a Gaussian process with independent and stationary increments, $W(0)=0$ and for all $0\leq s\leq t<\infty$, W(t)W(s)W(t)-W(s) is a Gaussian variable with mean zero and variance tst-s.
Let $\{N(t)\}_{t\geq 0}$ be a renewal process with iid inter-event times $\{J_{k}\}_{k\geq 1}$, independent of $\{X_{k}\}_{k\geq 1}$. Then $\{S_{N(t)}\}_{t\geq 0}$ is a time-changed random walk. If $\{N(t)\}_{t\geq 0}$ is a Poisson process with rate one and compounding rvs $\{X_{k}\}_{k\geq 1}$ are standard Gaussian then the scaled CTRW, $\{n^{-1/2}S_{N(nt)}\}_{t\geq 0}$, $n\geq 1$, converges weakly in appropriate Skorokhod topology to a BM (see [26]). Moreover, if $\{J_{k}\}_{k\geq 1}$ belong to the domain of attraction of a stable law with index $\alpha\in(0,1)$ and $\{X_{k}\}_{k\geq 1}$ belong to the domain of attraction of a standard Gaussian law, then the scaled CTRW, $\{n^{-\alpha/2}S_{N(nt)}\}_{t\geq 0}$, $n\geq 1$, $0<\alpha<1$, weakly converges in an appropriate Skorokhod topology to a time-changed Brownian motion (see [23, 26]), where the time changing component is an inverse stable subordinator (for definition see Section 2.2).
In non-commutative probability, the analogues of the Gaussian distribution and the Brownian motion are respectively the semi-circle distribution (also known as the free Gaussian distribution) and the (standard) free Brownian motion (FBM), say $\{B(t)\}_{t\geq 0}$ (see Section 2.3) for which the increments are freely independent, $B(t)=0$, and the variable B(t)B(s)B(t)-B(s) has a semi-circle distribution with variance tst-s for every $t\geq s$. At the same time, it is known that the algebraic limit variable of scaled Wigner matrices of increasing dimensions has the semi-circle distribution (see [33]), and that independent Wigner matrices are almost surely (a.s.) asymptotically free (see [1], [9]).
@@
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a classical probability space. Let $\mathcal{A}_{\mathbb{P}}=\cap_{1\leq p<\infty}L^{p}(\mathbb{P})$ be the algebra of rvs with all moments finite and let $\mathcal{M}_{n}(\mathcal{A}_{\mathbb{P}})$ be the algebra of $n\times n$ complex random matrices with entries from $\mathcal{A}_{\mathbb{P}}$. Then, the pair $(\mathcal{M}_{n}(\mathcal{A}_{\mathbb{P}}),\tau_{n})$ with $\tau_{n}=n^{-1}\mathbb{E}\mathrm{Trace}(\cdot)$ ($\mathbb{E}$ denotes expectation with respect to the probability $\mathbb{P}$) or $\tau_{n}=n^{-1}\mathrm{Trace}(\cdot)$ forms a non-commutative probability space (see Section 2.1 for a formal definition). Let $\{A_{n}(t)\}_{t\geq 0}$ be a collection from $\mathcal{M}_{n}(\mathcal{A}_{\mathbb{P}})$. Then, it is said to be algebraically (or jointly) convergent to a collection $\{A(t)\}_{t\geq 0}$ if, for any $p\in\mathbb{N}$ and distinct $t_{1},\dots,t_{p}\in[0,\infty)$, the matrices $A_{n}(t_{1}),\dots,A_{n}(t_{p})$ jointly converge to variables $A(t_{1}),\dots,A(t_{p})$ (see Section 2.1.2 for details).
Motivated by the above, we consider some sequences of patterned random matrices whose entries evolve as independent CTRWs. Let $\{X_{i,j,k}\}$ and $\{X_{i,k}\}$ be independent random variables with mean 0, variance $1$ and all moments uniformly bounded. For $i,j\geq 1$, let $S_{i,j,n}(t)=\sum_{k=1}^{[nt]}X_{i,j,k}$ and $S_{i,n}(t)=\sum^{[nt]}_{k=1}X_{i,k}$ $n\geq 1$ be independent random walks.
We construct patterned random matrices with these random walks as entries as follows. For either $d=1$ or $d=2$, let $L^{d}_{n}:\{1,2,\dots,n\}^{2}\rightarrow\mathbb{Z}^{d}$ (called link functions), with $L^{d}_{n+1}(i,j)=L^{d}_{n}(i,j)$ for all $1\leq i,j\leq n$. By abuse of notation, we denote these functions by $L^{d}$. This link is called symmetric if $L^{d}(i,j)=L^{d}(j,i)$ for all $i,j$. We consider matrices $A_{n}(t)=((S_{L^{d}(i,j),[nt]}))_{1\leq i,j\leq n}\}_{t\geq 0}$ in the NCP $(\mathcal{M}_{n}(\mathcal{A}_{\mathbb{P}}),\tau_{n})$. These matrices use $\{S_{i,n}\}$ or $\{S_{i,j,n}\}$ depending on whether $d=1$ or $d=2$. We also consider a time-changed version of these matrices (details given later). These matrices are symmetric if $L^{d}$ is symmetric.
@@
After recalling in Section 2, some preliminaries that we need, in Sections 3.1 and 3.2, we establish respectively, the marginal and joint convergence as defined above (that is, convergence of the expected average trace of any polynomial) of $\{n^{-1}A_{n}(t)\}_{t\geq 0}$ for suitable symmetric links $L$. These algebraic convergence results also yield the a.s. convergence of the empirical spectral distribution (ESD) of symmetric polynomials in $n^{-1}A_{n}(t),t_{1}<\cdots<t_{k}$ for every $\{t_{i}\}$ (and even independent copies of these).
For the choice of the Wigner link $L^{2}(i,j)=(\max(i,j),\min(i,j))$, $\{n^{-1}A_{n}(t)\}_{t\geq 0}$ converges algebraically to a free Brownian motion (FBM) $\{B(t)\}_{t\geq 0}$. The ESD of any symmetric polynomial in these matrices converges a.s. to the distributions determined by the moments of the corresponding (self-adjoint) polynomial of $\{B(t)\}_{t\geq 0}$.
@@
The choice of the link $L^{1}(i,j)=n/2-|n/2-|i-j||$, $1\leq i,j\leq n$, yields a Symmetric Circulant matrix. In this case, the matrices $\{n^{-1}A_{n}(t)\}_{t\geq 0}$ are commutative, and converge algebraically to $\{W(t)\}_{t\geq 0}$, whose finite dimensional moments agree with those of a BM. The ESD of any symmetric polynomial in these matrices converges to the distribution determined by the moments of the corresponding (self-adjoint) polynomial of $\{W(t)\}_{t\geq 0}$.
@@
In Section 3.3, with the above link function, we also consider the empirical process, $Y_{n}(t)=n^{-1}\sum_{r=1}^{n}\lambda_{r,n}(t)\mathbb{I}(U_{n}=r)$ where $\lambda_{r,n}(t),1\leq r\leq n$ are the eigenvalues of $n^{-1}A_{n}(t)$ and $U_{n}$ is a uniform random variable, taking values in $\{1,\ldots,n\}$, independent of all other variables. We show that $Y_{n}(\cdot)$ converges weakly to a BM in the Skorokhod space $D[0,\infty)$.
In Section 2.3 we introduce the time-changed FBM. In Section 4, we consider patterned random matrices with stopped (via fractional Poisson process) random walk entries as $\{A_{n}^{\alpha}(t)=((S_{L^{d}(i,j),N^{\alpha}(t)}))_{1\leq i,j\leq n}\}_{t\geq 0}$, $0<\alpha\leq 1$, where the time-changing component $\{N^{\alpha}(t)\}_{t\geq 0}$ is an independent fractional Poisson process (see Section 2.2 for the definition). We establish the algebraic convergence of the matrices $n^{-(1+\alpha)/2}A_{n}^{\alpha}(t)_{t\geq 0}$, for suitable symmetric links.
@@
For the link function $L^{2}(i,j)=(\max(i,j),\min(i,j))$, $\{n^{-(1+\alpha)/2)}A_{n}^{\alpha}(t)\}_{t\geq 0}$ converges algebraically to a time changed FBM $\{B(L^{\alpha}(t))\}_{t\geq 0}$, where $\{L^{\alpha}(t)\}_{t\geq 0}$ is an inverse stable subordinator independent of $B(\cdot)$. For any $t$, the moments of $B(L^{\alpha}(t))$ identify a unique probability distribution. Simulations (see Figure 1) show that its density function has a longer tail compared to the semi-circle density with variance $t$, and as $\alpha\to 1$, this density approaches the latter. As before, from the algebraic convergence, the a.s. convergence of the ESD of any symmetric polynomial can also be concluded.
Similarly, for the link $L^{1}(i,j)=n/2-|n/2-|i-j||$, $\{n^{-(1+\alpha)/2)}A_{n}^{\alpha}(t)\}_{t\geq 0}$ converges algebraically to a time changed BM $\{W(L^{\alpha}(t))\}_{t\geq 0}$, where $\{L^{\alpha}(t)\}_{t\geq 0}$ is an inverse stable subordinator independent of $W$. For any $t$, the moments of $W(L^{\alpha}(t))$ identify a unique probability distribution. As before, from the algebraic convergence, the a.s. convergence of the ESD of any symmetric polynomial can also be concluded.
In Section 4.1 we consider the link function L(i,j)=n/2|n/2|ij||L(i,j)=n/2-|n/2-|i-j||, $1\leq i,j\leq n$, and a process $Y_{n}^{\alpha}(t)=n^{-1}\sum_{r=1}^{n-1}\lambda_{r,n}^{\alpha}(t)\mathbb{I}(U_{n}=r)$, $t\geq 0$, where $\lambda_{r,n}(t)$ are the eigenvalues of $\{n^{-(1+\alpha)/2)}A_{n}^{\alpha}(t)\}_{t\geq 0}$. Analogous to the convergence results in Section 3.3, we show that this process converges weakly to $\{W(L^{\alpha}(t))\}_{t\geq 0}$ in Skorokhod topology, where $\{L^{\alpha}(t)\}_{t\geq 0}$ is an inverse stable subordinator that is independent of the Brownian motion $\{W(t)\}_{t\geq 0}$.
In Section 5 we deal with a few non-symmetric cases. The circulant, IID, and elliptic matrices produce algebraic limits as complex Brownian motion, free circulation motion, and free elliptic motion. We also discuss their time changed versions.
@@ 2.1 Non-commutative probability and (free) independence
A non-commutative probability space (NCP) is a pair $(\mathcal{A},\tau)$ with unital algebra $\mathcal{A}$ over $\mathbb{C}$ and a linear functional $\tau:\mathcal{A}\rightarrow\mathbb{C}$ such that $\tau(1_{\mathcal{A}})=1$, where $1_{\mathcal{A}}$ denotes the identity in $\mathcal{A}$. The state $\tau$ is called tracial if $\tau(ab)=\tau(ba)$ for all $a,b$ in $\mathcal{A}$. If $\mathcal{A}$ is a *-algebra, that is, there is an involution or *-operator $\mathcal{A}\rightarrow\mathcal{A}$, then a state $\tau$ on $\mathcal{A}$ is called positive if $\tau(a^{*}a)\geq 0$ for all $a\in\mathcal{A}$. The pair $(\mathcal{A},\tau)$ with *-algebra $\mathcal{A}$ and positive state $\tau$ is called *-probability space.